# Examples of robotic art Encountering Robotic Art: The Social, Material, and Temporal Processes of Creation with Machines

### v1

@@ 5.3. Process-Oriented Creation
As shown in our findings, the creation process itself embeds creative values and meanings, and experiencing the process can be pursued as the goal of creation with computing systems. For the robotic artists in our study, central artistic value was often placed on the creation process rather than the outcome. For example, in Alex’s robotic live drawing performance, the drawing process is more important than the drawn pattern on canvas. Techniques used, decisions made, or stimuli received by robots during creation or exhibition reflect artistic ideas and nuanced thinking, as seen in Sophie’s exploration of interactive decision-making in robotic drawing.
@@
The findings underscore the need to reconceptualize creation as encompassing more than just the process aimed at producing a final product; it also includes what we term post-creation. Distinct from repair or maintenance, post-creation entails anticipating and managing how an artifact evolves after its “completion” in the conventional sense. Specifically, we encourage creators to anticipate and strategically engage with the post-creation phase, considering potential changes to the artifact and their consequences for interactions with human users. For instance, during the creation process, creators may focus on possible material changes the artifact might undergo post-creation, allowing them to either mitigate or creatively exploit these potential changes. This expanded view of creation invites us to trace post-creation developments and to plan how our creative intentions can be embedded in its potential degradation, transformation, or evolution over time.
@@ Acknowledgements.
We are in debt to all robotic artists who participated in this study. We also thank the anonymous reviewers for their feedback. This study received no additional source of funding.References

### v2

@@ 5.3. Process-Oriented Creation
As shown in our findings, the creation process itself embeds creative values and meanings, and experiencing the process can be pursued as the goal of creation with computing systems. For the robotic artists in our study, artistic values were often placed on the creation process rather than the outcome. For example, in Alex’s robotic live drawing performance, the drawing process is more important than the drawn pattern on canvas. Techniques used, decisions made, or stimuli received by robots during creation or exhibition reflect artistic ideas and nuanced thinking, as seen in Sophie’s exploration of interactive decision-making in robotic drawing.
@@
The findings underscore the need to reconceptualize creation as encompassing more than just the process aimed at producing a final product; it also includes what we term post-creation. Distinct from repair, maintenance, or recycle, post-creation entails anticipating and managing how an artifact evolves after its “completion” in the conventional sense. Specifically, we encourage creators to anticipate and strategically engage with the post-creation phase, considering potential changes to the artifact and their consequences for interactions with human users. For instance, during the creation process, creators may focus on possible material changes the artifact might undergo post-creation, allowing them to either mitigate or creatively exploit these potential changes. This expanded view of creation invites us to trace post-creation developments and to plan how our creative intentions can be embedded in its potential degradation, transformation, or evolution over time.
@@ Acknowledgements.
We are in debt to all robotic artists who participated in this study. We also thank the anonymous reviewers for their feedback.References